Suppose that there is a .6 correlation between IQ (μ = 100, σ = 15) and verbal SAT score (μ = 500, σ = 100). What would be the variance of the estimate when predicting IQ from verbal SAT?

36

64

144

640
Answer #1

Variance of the estimate when predicting IQ from verbal SAT =Sy2*(1-r2) =15^2*(1-0.6^2)=144